# Capacity Gains in MIMO Systems with Few-Bit ADCs Using Nonlinear Analog Circuits

### v2

@@ Abstract
Analog to Digital Converters (ADCs) are a major contributor to the power consumption of multiple-input multiple-output (MIMO) receivers with large antenna arrays operating in the millimeter wave carrier frequencies. This is especially the case in large bandwidth communication systems, due to the sudden drop in energy-efficiency of ADCs as the sampling rate is increased above 100MHz. Two mitigating energy-efficient approaches which have received significant recent interest are i) to reduce the number of ADCs via analog and hybrid beamforming architectures, and ii) to reduce the resolution of the ADCs which in turn decreases power consumption. However, decreasing the number and resolution of ADCs leads to performance loss — in terms of achievable rates — due to increased quantization error. In this work, we study the application of practically implementable nonlinear analog operators such as envelope detectors and polynomial operators, prior to sampling and quantization at the ADCs, as a way to mitigate the aforementioned rate-loss. A receiver architecture consisting of linear analog combiners, nonlinear analog operators, and few-bit ADCs is designed. The fundamental information theoretic performance limits of the resulting communication system, in terms of achievable rates, are investigated under various assumptions on the set of implementable analog operators. Extensive numerical evaluations and simulations of the communication system are provided to compare the set of achievable rates under different architecture designs and parameters. Circuit simulations and measurement results, based on both 22 nm FDSOI CMOS technology and 65 nm Bulk CMOS transistor technologies, are provided to justify the power efficiency of the proposed receiver architecture deploying envelope detectors and polynomial operators.
@@ I Introduction
In order to accommodate the ever-growing demand for higher data-rates, the wireless spectrum has been continuously expanding over the past several decades. Particularly, millimeter wave (mm-wave) communication networks are being used in the fifth generation (5G) wireless systems to allow for larger channel bandwidths compared to earlier generation radio frequency (RF) systems which operate in frequencies below 6 GHz [3]. However, the energy consumption of components such as analog to digital converters (ADCs) increases significantly in mm-wave systems due to several factors as elaborated in the following. In theory, the power consumption of an ADC grows linearly with bandwidth, and the rate of increase is even more significant in practical implementations due to the excessive loss associated with the passive components at higher frequencies, which causes an abrupt drop in ADC energy-efficiency as the bandwidth is pushed past 100 MHz [4, 5, 6] For instance, the power consumption of current commercial high-speed ($\geq$ 20 GSample/s), high-resolution (e.g. 8-12 bits) ADCs is around 500 mW per ADC [7]. Furthermore, in order to mitigate the inherent high isotropic path loss and sensitivity to blockages at high frequencies, mm-wave systems must leverage directive narrow-beams, by using large antenna arrays to increase the antenna gain at both base stations (BS) and user-ends (UE) [8, 4, 9, 10]. For instance, 5G wireless networks envision hundreds of antennas at the BS and in excess of ten antennas at the UE [11].
@@
In conventional multiple-input multiple-output (MIMO) systems with digital beamforming, each antenna output is digitized separately by a dedicated ADC. This requires a large number of ADCs which are a significant source of power consumption in large bandwidth MIMO receivers [12, 13]. Analog and hybrid beamforming has been proposed as a way to mitigate ADC power consumption by reducing the number of ADCs. Under hybrid beamforming, the receiver terminals use a collection of analog beamformers to linearly combine the large number of received signals and feed them to a small set of ADCs. Additionally, in standard ADC design, power consumption is proportional to the number of quantization bins and hence grows exponentially in the number of output bits [14], which prohibits the use of high resolution ADCs.
@@
There has been extensive recent efforts to design receiver architectures and coding strategies using analog, hybrid, and digital beamforming with a small number of few-bit ADCs [15, 16, 17, 18, 12, 19, 20, 21, 22, 23, 24, 25, 26]. One method that has been proposed is to use digital beamforming architectures using one-bit ADCs [27]. A limitation of this approach is that digital beamforming requires an ADC per receiver antenna (two ADCs if input is considered in the complex domain), which leads to high power consumption. Furthermore, the restriction to one-bit ADCs may be prohibitively costly in terms of achievable rates. To mitigate this, hybrid beamforming has been proposed as an alternative approach. In hybrid beamforming, the receiver terminals in MIMO systems use a set of analog beamformers in the RF domain to combine the large number of analog signals at the receiver antennas and feed them to a small set of ADCs [18, 12, 19, 28]. Hybrid beamforming uses simple analog processing, linear processing in particular, to partially mitigate the rate loss due to low resolution quantization. In [29] a power-efficient hybrid MIMO receiver is presented where the analog and digital processing are jointly optimized by using task-specific quantization techniques. The hybrid beamforming approach was further extended in [30], where the use of hybrid blockwise architectures, consisting of delay elements, was considered. This allows for temporal linear processing of the received signals in the analog domain, providing additional degrees of freedom in choosing the processing function. Additionally, adaptive thresholds ADCs, were considered in [30], which allow for modifying the ADC thresholds based on past quantization outputs. This has similarities with the operation of successive approximation register (SAR) ADCs [31, 32]. The latter two approaches improve the channel capacity compared with that of the conventional hybrid beamforming architectures, however, the practical implementation of high precision analog delay elements, and SAR ADCs with high sampling rates is challenging due to synchronization issues and delay accuracy limitations.
@@
In this work, we argue that the aforementioned digital, hybrid, and analog beamforming approaches suffer from the curse of low dimensions phenomenon, elaborated on in Section III. We further quantify the rate-loss due to this phenomenon, and provide solutions to mitigate it. Particularly, we consider the use of nonlinear analog operators — such as envelope detectors and low degree polynomial operators — prior to sampling and quantization, as a way to mitigate the rate-loss due to coarse quantization and increase channel capacity of MIMO systems with few-bit ADCs. The power consumption and circuit design for receiver architectures deploying these nonlinear operators based on measurement and simulation results in 22 nm and 65 nm CMOS technologies are provided in Section VIII.
@@
  • To characterize the channel capacity under analog beamforming when envelope detectors are used for analog signal processing. The capacity expression is derived in terms of the power budget for analog processing components and ADCs, which in turn determines the number of ADCs, $n_{q}$, number of output levels of each ADC, $\ell$, and number of concatenated envelope detectors, $\delta_{env}$ (Theorems 1-3).
@@
  • To characterize the high SNR capacity under analog beamforming when polynomial operators are used for analog signal processing. The high SNR capacity and inner-bounds to the low SNR achievable rates are provided in terms of the power budget of analog processing components and the ADCs, which in turn determine the number of ADCs, $n_{q}$, number of output levels of each ADC, $\ell$, and the polynomial degree, $\delta_{poly}$ (Theorem 4).
  • To provide a receiver architecture for hybrid beamforming using envelope detectors for high data rate communication with QAM demodulation (Theorem 5).
  • To provide computational methods for finding the set of achievable rates and quantifying the gains due to nonlinear analog processing under the proposed analog and hybrid beamforming architectures, and to provide explanations of how these gains change as SNR, $n_{q}$, $\ell$, and $\delta$ are changed.
@@
  • To provide circuit designs and associated performance simulations for implementing polynomials of degree up to four and concatenated envelope detector sequences with a pair of envelope detectors, and to evaluate their power consumption.
The organization of the remainder of the paper is as follows: In Section II, we provide an overview of prior works under both digital and hybrid beamforming and elaborate on the motivation behind considering nonlinear processing in the analog domain. We provide several motivating examples to illustrate the curse of low dimensions phenomenon and motivate the use of non-linear analog operators in Section III. In Section IV, we consider the scenario where the set of implementable analog processing functions consists of those generated by collections of envelope detectors, and propose receiver architectures, design coding strategies, and derive the fundamental performance limits in terms of achievable rates. In Section V, we study receiver architectures and coding strategies associated with analog polynomial operators. In Section VI, we consider a hybrid beamforming architecture that is equipped with envelope detectors. We provide achievable rates under this architecture in a high SNR regime. A numerical analysis for channel capacity with respect to each of the described architectures is provided in Section VII. Here we characterize the gains that result from the use of nonlinear processing in the analog domain. Transceiver Circuit designs and simulations evaluating the power consumption of the nonlinear processing operators including sequences of concatenated envelope detectors and polynomial operators are provided in Section VIII.
@@
Notation: Upper-case letters such as $X,\mathbf{Y},\mathbf{H}$ are used to represent random variables, vectors, and matrices, and lower-case letters such as $x,\mathbf{y},\mathbf{h}$ represent their realizations. The set $\{1,2,\cdots,n\},n\in\mathbb{N}$ is represented by $[n]$. The vector $(x_{1},x_{2},\ldots,x_{n})$ is written as $x(1\!\!:\!\!n)$ and $x^{n}$, interchangeably, and $(x_{k},x_{k+1},\cdots,x_{n})$ is denoted by $x(k:n)$. The $i$th element is written as $x(i)$ and $x_{i}$, interchangeably. We write $||\cdot||_{2}$ to denote the $L_{2}$-norm. An $n\times m$ matrix is written as $h(1\!\!:\!\!n,1\!\!:\!\!m)=[h_{i,j}]_{i,j\in[n]\times[m]}$, , its $i$th column is $h(:,i),i\in[m]$, and its $j$th row is $h(j,:),j\in[m]$. We write $\mathbf{x}$ and $\mathbf{h}$ instead of $x(1\!\!:\!\!n)$ and $h(1\!\!:\!\!n,1\!\!:\!\!m)$, respectively, when the dimension is clear from context. The notation $\mathbf{x}^{H}$ is used to denote the hermitian of $\mathbf{x}$. Sets are denoted by calligraphic letters such as $\mathcal{X}$, families of sets by sans-serif letters such as $\mathsf{X}$, and collections of families of sets by $\mathscr{X}$.
II Preliminaries
II-A The Classical Information Theoretic MIMO Communications Model
The classical model for Point-to-Point (PtP) MIMO communications is shown in Figure 1(a). The model has been considered under various scenarios such as availability of channel state information (CSI) at one or both terminals, time-varying channels, and lossy transmission of sources over MIMO channels, e.g., [33, 34]. A consequential idealization in the classical model is that no restrictions are imposed on the choice of encoding and decoding functions $g(\cdot)$ and $f(\cdot)$ at the transmitter and receiver terminals, respectively. At the receiver side, the justification is that the decoding function may be implemented by first constructing a faithful digital representation of the analog signal using high resolution ADCs and then performing a blockwise decoding operation on these signals in the digital domain.
Figure 1: MIMO system $n_{t}$ transmit antennas, $n_{r}$ receive antennas, and $n_{q}$ ADCs. The receiver model (dashed box) consists of analog processing $f_{a}(\cdot)$, analog to digital conversion (ADC), and digital processing $f_{d}(\cdot)$ modules.
However, this idealistic model does not capture the limitations on the resolution of the ADCs and the limitations of analog circuit design in mm-wave systems mentioned in the introduction. In this work, we focus on the receiver side, and consider a system model which captures the aforementioned practical considerations in the receiver architecture.
We consider a MIMO channel whose input and output [^1] To simplify notation, we have considered real-valued variables. The derivations can be extended to complex variables in a straightforward manner. $(\mathbf{X},\mathbf{Y})\in\mathbb{R}^{n_{t}}\times\mathbb{R}^{n_{r}}$ are related through $\mathbf{Y}=\mathbf{h}\mathbf{X}+\mathbf{N}$, where, $\mathbf{h}\in\mathbb{R}^{n_{r}\times n_{t}}$ is the (fixed) channel matrix, $n_{t}$ is the number of transmit antennas, $n_{r}$ is the number of receive antennas, and $\mathbf{N}\in\mathbb{R}^{n_{r}}$ is a jointly Gaussian noise vector. The channel input has average power constraint $P_{T}$, i.e. $\mathbb{E}(\|\mathbf{X}\|_{2}^{2})\leq P_{T}$, where $\|\cdot\|_{2}$ denotes the $\ell_{2}$ distance. The receiver model is shown in Figure 1. The model encompasses the analog beamforming, hybrid beamforming, and digital beamforming models used in prior works as well as the beamforming models introduced in the subsequent sections. As observed from the figure, the choice of the decoding function at the receiver is restricted by limitations on the number of ADCs, $n_{q}\in\mathbb{N}$ — which is in turn restricted by the power budget at the receiver side — and the set of implementable nonlinear analog functions $f_{a}(\cdot)$. In general, we assume that the analog function $f_{a}(\cdot)$ is chosen from a predefined set of implementable functions $\mathcal{F}_{a}$. Each of these restrictions are elaborated on in the subsequent sections.
Figure 2: The receiver architecture consists of an analog/hybrid beamformoing module, an elementwise analog operator $f_{a}(\cdot)$, $n_{q}$ low-resolution ADCs, and a blockwise digital operator $f_{d}(\cdot)$ with blocklength $n$. $Y$ represents the received signal, $\widehat{M}$ is the message reconstruction, and $[\Theta]$ is the message set.
Formally, the receiver (Figure 2), consists of:
i) A beamforming module characterized by $\mathbf{V}\in\mathbb{R}^{n_{s}\times n_{r}}$ which takes $\mathbf{Y}\in\mathbb{R}^{n_{r}}$ as input and outputs $\mathbf{\widetilde{Y}}\in\mathbb{R}^{n_{s}}=\mathbf{V}\mathbf{Y}$ using a collection of phase shifters.
ii) A set of elementwise analog processing functions $f_{j}\in\mathcal{F}_{a},j\in[n_{q}],\delta\in\mathbb{N}$ operating on the beamformer output $\mathbf{Y}$ to produce the vector $W(1:n_{q})$, where $W(j)=f_{j}(\mathbf{Y}),j\in[n_{q}]$. Note that both $\mathbf{V}$ and $f_{a}(\cdot)$ capture operations performed in the analog processing of the received signals. Although the two operators are implemented in different frequency bands — the carrier frequency and baseband, respectively — they can be effectively combined together into one module for the purpose of the derivations considered in this work. Consequently, in the following, we represent the analog processing module by $f_{a}(\cdot)$, and take $\mathbf{V}$ as the identity matrix for conciseness. We consider two classes of analog linear processing functions:
Concatenated Sequences of Envelope Detectors: In Section IV, we consider the use of concatenated sequences of envelope detectors for analog signal processing. In this context, we define the set of implementable nonlinear analog functions as:
$\mathcal{F}^{\delta}_{env}=\{f(y)=A_{s}(y,a^{s}),y\in\mathbb{R}|s\in[\delta],a^{s}\in\mathbb{R}^{s}\},\quad\delta\in\mathbb{N},$
where $A_{1}(y,a)\triangleq|y-a|,y,a\in\mathbb{R}$ and $A_{s}(y,a^{s})\triangleq A_{1}(A_{s-1}(y,a^{s-1}),a_{s})=|A_{s-1}(y,a^{s-1})-a_{s}|,s\in\mathbb{N}$. That is, $\mathcal{F}^{\delta}_{env}$ consists of all functions which can be generated using sequences of $s\leq\delta$ concatenated envelope detectors with thresholds $a_{1},a_{2},\cdots,a_{s}$, respectively. The restriction to a limited number of envelope detectors is due to limitations in analog circuit design, and the implementability of such functions is justified by the circuit designs and simulations provided in Section VIII.
Analog Polynomial Operators: In Section V, we consider the use of polynomial operators. That is, we limit the set of implementable nonlinar analog functions to:
$\mathcal{F}^{\delta}_{poly}\triangleq\{f(y)=\sum_{i=0}^{\delta}{a_{i}}y^{i},y\in\mathbb{R}|a_{i}\in\mathbb{R},i=0,1,\cdots,\delta\},\quad\delta\in\mathbb{N},$
In this case, $\mathcal{F}^{\delta}_{poly}$ consists of all polynomial functions of degree at most $\delta$, where the thresholds are determined by the roots of the polynomials. Similar to the case with envelope detectors, here the limitation on the degree of the polynomial functions is determined by limitations in the corresponding circuits design and is studied in Section VIII.
iii) A set of ADCs, whose number $n_{q}$ and resolution [^2] Generally, one could assume different resolutions for each of the ADCs, however, we assume all ADCs have the same resolution to simplify the circuit design and implementation. $\ell$ is determined by a given power budget $P_{ADC}$. To elaborate, theoretically, the power consumption of an ADC is given by $CVf_{s}2^{n}$, where $C$ is the capacitive load, $V$ is the voltage, $f_{s}$ is the sampling frequency, and $n$ is the number of output bits [35]. In our model, we define $\alpha\triangleq CVf_{s}$. As a result, if a one-bit ADC has a unit power consumption of $\alpha$, then an $n$-bit ADC has $2^{n}\alpha$ power consumption. Consequently, $n_{q}$ and $\ell$ must belong to the set
$\mathcal{N}(P_{ADC})\triangleq\{(n_{q},\ell)|\alpha n_{q}{\ell}\leq P_{ADC}\}.$
The $n_{q}$ ADCs have threshold vectors $t(j,1:\ell-1)\in\mathbb{R}^{\ell-1},j\in[n_{q}]$, operating on the vector $W(1:n_{q})$ and producing $\widehat{W}(1:n_{q})$, where
$\widehat{W}(j)=k\quad\text{ if }\quad W(j)\in[t(j,k),t(j,k+1)],k\in[0,\ell-1],$
where $j\in[n_{q}]$ and we have defined $t(j,0)\triangleq-\infty$ and $t(j,\ell)\triangleq\infty$. We call $t(1\!\!:\!n_{q},1\!\!:\!\ell-1)$ the threshold matrix. It is assumed that $0<t(i,j)<t(i,j^{\prime}),i\in[n_{q}],j,j^{\prime}\in[\ell-1],j<j^{\prime}$ [^3] Note that the assumption $0<t(i,j)$ does not loose generality since $0\leq|y|,\forall y\in\mathbb{R}$. Hence, a negative threshold would yield trivial ADC output..
iv) A digital processing module represented by $f_{d}:\{0,1,\cdots,\ell-1\}^{n\times n_{q}}\to[\Theta]$, operating on the block of ADC outputs after $n$-channel uses $\widehat{W}(1\!\!:\!\!n,1\!\!:\!\!n_{q})$. After the $n$th channel-use, the digital processing module produces the message reconstruction $\widehat{M}=f_{d}(\widehat{W}(1\!:\!n,1\!:\!n_{q}))$. The communication system is characterized by $(P_{T},P_{ADC},{h},\delta)$, and the transmission system by $(n,n_{q},\ell,\Theta,e,f^{n_{q}},t(1\!:\!n_{q},1\!:\!\ell-1),f_{d})$, where $(n_{q},\ell)\in\mathcal{N}(P_{ADC})$, $f^{n_{q}}=(f_{1},f_{2},\cdots,f_{n_{q}})$ and $f_{j}\in\mathcal{F}_{env}^{\delta},j\in[n_{q}]$ , and $e(\cdot)$ is such that the channel input satisfies the average power constraint. Achievability and probability of error are defined in the standard sense. The capacity maximized over all implementable analog functions is denoted by $C_{env}(P_{T},P_{ADC},h,\delta)$, when using sequences of envelope detectors and by $C_{poly}(P_{T},P_{ADC},h,\delta)$, when using polynomial operators.
II-B Summary of Prior Works
In the following, we provide a brief summary of prior works on digital and hybrid beamforming, and explain how each scenario can be encompassed as a special case of the formulation described in the prequel.
Prior Works on Digital Beamforming: In digital beamforming, each antenna is directly connected to its dedicated ADC, i.e., $n_{r}=n_{q}$ and $f_{a}(\cdot)$ is the identity function [36, 37, 38, 39, 40]. In [37], an approximate expression for the high signal-to-noise ratio (SNR) capacity of the single-input multiple-output (SIMO) system with digital beamforming was derived when one-bit ADCs are used; it was shown that the capacity is well-approximated by $\log{(4n_{q}+1)}$. Note that ideally, one would expect that at high SNRs, when the channel is almost noiseless, $n_{q}$ bits of information can be recovered per channel-use using $n_{q}$ one-bit ADCs. It can be observed that there is a significant gap between the performance of the digital beamforming architecture ($\log{(4n_{q}+1)}$ bits/channel-use) and the ideal performance ($n_{q}$ bits/channel-use). In this work, we argue that this gap is due to what we call the ‘curse of low dimensions’, further discussed in Section III and provide solutions to breach this performance gap.
Prior Works on Hybrid Beamforming: Hybrid beamforming architectures use a set of analog
Figure 3: The figure shows the hybrid beamforming architecture studied in prior works. The matrix $\mathbf{V}^{n_{r}\times n_{q}}$ characterizes the analog linear combiner, and $t^{n_{q}}$ is the vector of ADC thresholds.
beamformers to linearly combine the received signals in the analog domain [41, 42, 19], i.e., $f_{a}(\cdot)$ is a linear function which can be captured by matrix $\mathbf{V}^{n_{q}\times n_{r}}$ (Figure 3). This linear analog signal processing improves performance, in terms of achievable rates, by rotating the received signal such that the information-loss in the quantization step is reduced. For Gaussian SISO channels, hybrid beamforming with non-zero threshold ADCs was investigated in [43]. In MIMO communications, the use of analog linear combiners to reduce the quantization rate-loss was investigated in [19, 44]. Similar to the digital beamforming scenario alluded to above, the high SNR achievable rate in systems with hybrid beamforming has been shown to grow logarithmically in $n_{q}$ [44]. This is significantly lower than the ideal high SNR performance of $n_{q}$ bits/channel-use. Similar to digital beamforming, we argue that this loss in performance is due to the ‘curse of low dimensions’ and provide solutions to improve achievable rates.

### v3

# Non-Linear Analog Processing in MIMO Systems with Coarse Quantization

@@ Abstract
Analog to digital converters (ADCs) are a major contributor to the power consumption of multiple-input multiple-output (MIMO) receivers in large bandwidth millimeter-wave systems. Prior works have considered two mitigating solutions to reduce the ADC power consumption: i) decreasing the number of ADCs via analog and hybrid beamforming, and ii) decreasing the ADC resolution, i.e., utilizing one-bit and few-bit ADCs. These mitigating solutions lead to performance loss in terms of achievable rates due to increased quantization error. In this work, the use of nonlinear analog operators such as envelope detectors and polynomial operators, prior to sampling and quantization is considered, as a way to reduce the aforementioned rate-loss. The receiver architecture consists of linear combiners, nonlinear analog operators, and few-bit ADCs. The fundamental performance limits of the resulting communication system, in terms of achievable rates, are investigated under various assumptions on the set of implementable analog operators. Extensive numerical evaluations are provided to evaluate the set of achievable rates and the power consumption of the proposed receiver architectures. Circuit simulations and measurement results, based on both 22 nm FDSOI CMOS technology and 65 nm Bulk CMOS transistor technologies, are provided to justify the power efficiency of the proposed receiver architectures.
@@ I Introduction
In order to accommodate the demand for higher data-rates, the wireless spectrum has been continuously expanding over the past several decades. The millimeter-wave (mm-wave) spectrum is utilized in the fifth generation (5G) wireless systems to allow for larger channel bandwidths compared to earlier generation radio frequency systems, which operate in frequencies below 6 GHz [3]. The energy consumption of components such as analog to digital converters (ADCs) increases significantly in mm-wave systems. In theory, the power consumption of an ADC grows linearly with bandwidth, and the rate of increase is even more significant in practice due to the excessive loss associated with the passive components at higher frequencies, which causes an abrupt drop in ADC energy-efficiency as the bandwidth is pushed past 100 MHz [4]. For instance, the power consumption of commercial high-speed ($\geq$ 20 GSample/s), high-resolution (8-12 bits) ADCs is around 500 mW per ADC [5].
@@
The ADC power consumption is particularly prohibitive in mm-wave systems utilizing large antenna arrays and digital beamforming archiectures. To elaborate, in order to mitigate the inherent high isotropic path loss at high frequencies, mm-wave systems utilize directive narrow-beams generated by large antenna arrays [6]. For instance, 5G wireless networks envision hundreds of antennas at the base station (BS) and in excess of ten antennas at the user-end (UE) [7]. In conventional multiple-input multiple-output (MIMO) systems with digital beamforming, each antenna output is digitized separately by a dedicated ADC. This requires a large number of ADCs which are a significant source of power consumption in large bandwidth MIMO receivers [8, 9]. Analog and hybrid beamforming have been proposed to mitigate the ADC power consumption by reducing the number of ADCs. In hybrid beamforming, the receiver uses a set of analog beamformers in the radio frequency (RF) domain to combine the large number of analog signals at the receiver antennas and feed them to a small set of ADCs [10, 8, 11]. In analog beamforming, a single beamformer is used to linearly combine the received signals [12, 13].
@@
In addition to these beamforming architectures, an alternative approach to reduce ADC power consumption is to reduce the resolution of the ADCs. To elaborate, in the standard ADC design, power consumption is proportional to the number of quantization bins and hence grows exponentially in the number of output bits [14], which prohibits the use of high resolution ADCs. There has been extensive recent efforts to design receiver architectures and coding strategies using analog, hybrid, and digital beamforming with few-bit ADCs [15, 13, 16, 10, 8, 17, 18, 19, 20, 21]. The restriction to few-bit ADCs leads to reduced communication rates compared to when high resolution ADCs are utilized. In hybrid beamforming with few-bit ADCs, simple analog processing, linear processing in particular, is used to partially mitigate the rate loss due to low resolution quantization [22, 23, 20, 24, 25]. In [26], a power-efficient hybrid MIMO receiver is presented where the analog and digital processing are jointly optimized by using task-specific quantization techniques. The hybrid beamforming approach was further extended in [27], where hybrid blockwise architectures, consisting of delay elements, were considered. This allows for temporal linear processing of signals in the analog domain, providing additional degrees of freedom in choosing the processing function. Adaptive thresholds ADCs were considered in [27], which allow for modifying the ADC thresholds based on past quantization outputs. This method has similarities with successive approximation register (SAR) ADCs [28]. The latter two approaches improve the channel capacity compared with conventional hybrid beamforming architectures, however, the practical implementation of high precision analog delay elements and SAR ADCs with high sampling rates is challenging due to synchronization issues and delay accuracy limitations.
@@
In this work, we argue that digital, hybrid, and analog beamforming approaches suffer from a phenomenon which we call the curse of low dimensions (Section III). We quantify the rate-loss due to this phenomenon, and provide solutions to mitigate it. Particularly, we consider nonlinear analog operators — such as envelope detectors and low degree polynomial operators — prior to sampling and quantization, as a way to mitigate the rate-loss due to coarse quantization. The power consumption and circuit design for receiver architectures deploying these nonlinear operators based on measurements and simulations in 22 nm and 65 nm CMOS technologies are provided in Section VIII.
@@
  • To characterize the channel capacity under analog beamforming when envelope detectors are used for analog signal processing. (Theorems 1-3).
  • To characterize the capacity under analog beamforming when polynomial operators are used. (Theorem 4).
@@
  • To introduce a receiver architecture for hybrid beamforming using envelope detectors, and to provide the high signal-to-noise ratio (SNR) capacity and inner-bounds to the low SNR achievable rates in terms of the total ADC power budget $P_{ADC}$, number of output levels of each ADC, $\ell$, and the polynomial degree, $\delta_{poly}$ (Theorem 5).
@@
  • To provide circuit designs and associated performance simulations for implementing polynomials of degree up to four and pairs of concatenated envelope detectors, and to evaluate their power consumption. (Sections VII and VIII)
  • To evaluate the rate-power tradeoff under linear and nonlinear analog processing using the rate derivations in our theoretical analysis and the power figures in our circuit simulations. (Sections VII and VIII)
@@
Notation: Sets are denoted by calligraphic letters such as $\mathcal{X}$, families of sets by sans-serif letters such as $\mathsf{X}$. The set $\{1,2,\cdots,n\}$ is represented by $[n]$. $\mathcal{X}^{c}$ denotes the complement of $\mathcal{X}$. The vector $(x_{1},x_{2},\ldots,x_{n})$ is written as $x^{n}$, and $(x_{k},x_{k+1},\cdots,x_{n})$ is denoted by $x_{k}^{n}$. The $i$th element of $x^{n}$ is written as $x_{i}$. An $n\times m$ matrix is written as $h^{n\times m}=[h_{i,j}]_{i\in[n],j\in[m]}$, its $i$th row is denoted by $h_{i,:}=[h_{i,j}]_{j\in[m]}$ and its $j$th column is $h_{:,j}=[h_{i,j}]_{i\in[n]}$. The $n\times n$ identity matrix is denoted by $\mathbf{I}_{n}$. We use bold-face letters such as $\mathbf{x}$ and $\mathbf{h}$ instead of $x^{n}$ and $h^{n\times m}$, respectively, to represent vectors and matrices when the dimension is clear from the context. $\mathbf{x}^{H}$ denotes the hermitian of $\mathbf{x}$. We write $||\cdot||_{2}$ to denote the $L_{2}$-norm. Upper-case letters such as $X$ represent random variables, and lower-case letters such as $x$ represent their realizations. Similarly, random vectors and random matrices are denoted by upper-case letters such as $\mathbf{X}$ and $\mathbf{H}$, respectively. For a Gaussian random vector $\mathbf{X}$ with mean vector $\boldsymbol{\mu}$ and covariance matrix $\mathbf{\Sigma}$, we write $\mathbf{X}\sim\mathcal{N}(\boldsymbol{\mu},\mathbf{\Sigma})$.
II Problem Formulation
Figure 1: (Top) MIMO communication setup consisting of message $M$, channel input $\mathbf{X}$, channel output $\mathbf{Y}$, and message reconstruction $\widehat{M}$. (Bottom) receiver consisting of an analog processing module $\mathbf{f}_{a}(\cdot)$, ADC input $\mathbf{W}$, ADC module $\mathbf{a}(\cdot)$, ADC output $\mathbf{\widehat{W}}$, and a digital processing module $\mathbf{f}_{d}(\cdot)$.
II-A System Model
We consider a MIMO channel with $n_{t}$ transmit antennas and $n_{r}$ receive antennas. The input and output [^1] To simplify notation, we have considered real-valued variables. The derivations can be extended to complex variables in a straightforward manner. $(\mathbf{X},\mathbf{Y})\in\mathbb{R}^{n_{t}}\times\mathbb{R}^{n_{r}}$ are related through $\mathbf{Y}=\mathbf{h}\mathbf{X}+\mathbf{N}$, where $\mathbf{h}\in\mathbb{R}^{n_{r}\times n_{t}}$ is the (fixed) channel gain matrix and $\mathbf{N}\sim\mathcal{N}(0,I_{n_{r}})$ is a vector of Gaussian variables with zero mean and unit variance. The channel input has average power constraint $P_{T}$, i.e., $\mathbb{E}(\|\mathbf{X}\|_{2}^{2})\leq P_{T}$.
The receiver model is shown in the bottom figure in Figure 1. In general, upon receiving the channel output $\mathbf{Y}$, the receiver may perform analog processing operations. For instance, in hybrid beamforming the receiver applies an affine transformation on the received signal using an analog linear combiner, and in analog beamforming, phase shifters are used to produce a linear combination of the received signals. The analog processing operations are represented by the function $\mathbf{f}_{a}(\cdot)$. The output of $\mathbf{f}_{a}(\cdot)$ has dimension equal to the number of ADCs available at the receiver. That is, the function $\mathbf{f}_{a}(\cdot)$ takes the $n_{r}$-dimensional vector $\mathbf{Y}$ as input, and outputs an $n_{q}$-dimensional vector $\mathbf{W}=\mathbf{f}_{a}(\mathbf{Y})$, where $n_{q}$ is the number of ADCs. We denote $\mathbf{f}_{a}=(f_{1},f_{2},\cdots,f_{n_{q}})$, where $f_{j}:\mathbb{R}^{n_{r}}\to\mathbb{R},j\in[n_{q}]$. The analog vector $\mathbf{W}$ is fed to the ADCs to produce the digitized signal $\mathbf{\widehat{W}}$. The operation of the ADCs is described in Section II-B. The receiver then performs blockwise digital processing on the ADC outputs to produce the reconstruction $\widehat{M}$ of the transmitted message $M$. The digital processing operations performed to decode the message are represented by $\mathbf{f}_{d}(\cdot)$.
II-B The ADC Module
We assume that the receiver is equipped with a collection of ADCs. To simplify the notation, we assume that each of the ADCs have $\ell$ output levels, i.e., the ADCs have the same resolution. An ADC is modeled as a mapping from a continuous-valued input $y$ to a discrete-valued output $\hat{y}$. For an ADC with $\ell$ output levels, we define a threshold vector $t^{\ell-1}=(t_{1},t_{2},\cdots,t_{\ell-1})$. The output is then given by:
$\hat{y}=\begin{cases}0\qquad&\text{if }\quad y<t_{1}\\ i&\text{if }\quad\exists i\in[\ell-2],t_{i}\leq y<t_{i+1}\\ \ell-1&\text{if }\quad t_{\ell-1}\leq y\end{cases},$ (1)
where $t_{1}<t_{2}<\cdots<t_{\ell-1}$. For a receiver equipped with $n_{q}$ ADCs each with $\ell$ output levels, the threshold matrix is defined as $t^{n_{q}\times{(\ell-1)}}$, where $t_{i,:},i\in[n_{q}]$ is the threshold vector corresponding to the $i$th ADC.
Theoretically, the power consumption of an ADC is equal to the product of its capacitive load C, voltage V, sampling frequency $\omega_{s}$, and number of output levels $\ell$ [29]. That is, the power consumption of a single $\ell$-level ADC is $\alpha\ell$, where we have defined $\alpha\triangleq CV\omega_{s}$. Given a total ADC power budget $P_{ADC}$ and number of output levels $\ell$, the maximum number of ADCs at the receiver is given by $n_{q}=\lfloor\frac{P_{ADC}}{\alpha\ell}\rfloor$.
II-C Conventional Beamforming Architectures